# Prosody Analysis of Audiobooks

### v2

@@ 5.5.2 Sequential Prediction
Table 6: Test MSE for joint prediction of prosody attributes with LSTM
Table 6 shows the test MSE for the three attributes. The LSTM model trained on MPNet phrase embeddings, with sequence length 3 shows the best performance. We use this model further downstream for text-to-speech audio generation and human evaluation.
5.5.3 Prediction on Character Dialogue
@@
We further examine this effect by looking at the prosody attribute prediction on dialogues only. Table 7 shows the MSE on a subset of the test data containing only dialogues. Pitch and volume attributes on dialogues are predicted with lower error compared to all input.
@@
Table 7: Test MSE for joint prediction of prosody attributes on dialogues
Character Embeddings:
More subtle information about the nature of individual characters (e.g. hero vs. villain, old vs. young) inform actor’s readings. We conducted an experiment to augment our models by concatenating 908d character embeddings (Inoue et al., 2022) to text representations before applying PCA for dimension reduction.
@@
Table 8: Test MSE for joint prediction of prosody attributes on dialogues using character embeddings (CE) different PCA dimensions.
Table 8 shows the MSE on plain audio generated without SSML attributes and enhanced audio generated with SSML attributes predicted from text and character embeddings. The results improve over our baseline, but not not over our best models.
6 Generating and Evaluating Text-to-Speech Audiobooks
We use the best performing model (LSTM with sequence length 3, and MPNet embeddings as input) to generate predictions for each book in the test set. We use a sliding window for inference, and compute the final prediction for each text input as the mean of its prediction in all windows.
To convert the predicted z-scores into SSML attributes, we compute the mean and standard deviation for the pitch and volume in each human-read audio file in the test set, and use these values to convert the prosody attributes into absolute values (Hz for pitch and dB for volume). Using the mean pitch/volume as reference, we convert these into relative values as required by the SSML specification (as described in Section 2.1). For rate of speech, we use a fixed value for mean and standard deviation (100 and 50 respectively) and perform the same computation. Note that we use these values as we have a reference point from the human-read audios, but these can be adapted as per the TTS voice to introduce more prosodic variations.
6.1 Human Evaluation
We have demonstrated that our pitch, volume, and rate models successfully predict the prosody of human audiobook readers. However, it is not apparent whether human listeners will prefer text-to-speech synthesis augmented by our analysis over SSML models developed by commercial TTS vendors. This is a very difficult baseline for us to try to beat, given the vast training resources deployed to build commercial TTS systems.
We performed a Mechanical Turk study to evaluate our models. For each of the 24 books in the test set, we identify the phrases with the highest absolute values of predictions across the three prosody attributes. We then generate audio for this phrase text along with one context phrase before and after, using the Google Cloud Text-to-Speech API  [^8] https://cloud.google.com/text-to-speech with the voice en-US-Standard-A. We produce two different audio variants for each text:
@@ 6.1 Human Evaluation
  • SSML: We enclose each phrase in a <prosody> tag with the three predicted attributes specified: see example in Figure 2.
For each pair of audios, we ask five annotators on Amazon Mechanical Turk to rate which audio sounds better and more natural with intonations similar to human readers.
SSML Plain Neutral Count
5 0 0 1
4 1 0 2
3 1 1 2
3 2 0 7
2 2 1 2
2 3 0 5
1 3 1 1
1 4 0 2
0 5 0 2
Table 9: Annotator preference counts: denotes whether majority annotators preferred SSML or Plain text.
Table 9 shows the results of the evaluation. Each row shows the annotator preferences, and the number of books from the test set with this distribution. For example, row 2 means that there are 2 pairs of audios for which 4 annotators preferred the SSML-enhanced version over the plain text version. For 12 out of 24 test sample pairs, a majority of the annotators preferred the SSML-enhanced version.
Qualitative Analysis:
7 out of the 10 pairs in which the majority preferred plain text had phrases with higher values for pitch and volume in the SSML-enhanced version. Also, for 8 of 12 pairs in which the majority preferred the SSML version, the rate of speech was lower than the default.
7 Conclusion
We have developed an aligned text-audio book corpus, and deployed it to study both reader behavior and improve prosody models for TTS generation of audiobooks. We have shown that actors customize voices to represent specific characters, and adopt more expressive voices than those generated by conventional TTS systems, motivating the development of full-book analysis methods to understand/predict a reader’s goals in narrating a text.
Limitations
  • Our approach produces expressive audio through a sequence-to-sequence model producing annotated SSML to feed to commercial TTS system, instead of directly-trained neural text-to-speech models (text in, audio out). Presumably higher performance scores would follow from the more direct approach, given enough training data.
  • Our sequence-to-sequence approach makes it difficult to control expressiveness at a low level of granularity: word-level prosody cannot be properly controlled because the generated speech sounds disconnected. We believe this is why our models did not perform better at human evaluation.
  • Our research relies on paid access to a commercial TTS system, which limits the size of our experiments. Available open source TTS systems produced significantly poorer speech, make it less attractive as a foundation for our work.

### v3

@@ V-C Sequential Prediction
TABLE II: Test MSE for contextual joint prediction of prosody attributes with LSTM
Table II shows the test MSE for the three attributes. The LSTM model trained on MPNet phrase embeddings, with sequence length 3 shows the best performance. We use this model further downstream for text-to-speech audio generation and human evaluation.
V-D Prediction on Character Dialogue
@@
We further examine this effect by looking at the prosody attribute prediction on dialogues only. Table III shows the MSE on a subset of the test data containing only dialogues. Predicting pitch and volume attributes on dialogues has a lower error compared to all input.
@@
TABLE III: Test MSE for contextual joint prediction of prosody attributes on dialogues
V-E Character Embeddings
To include more subtle information about the nature of individual characters (e.g. hero vs. villain, old vs. young) that may inform actors’ readings, we appended 908-dimensional character embeddings [12] to the input text representation and applied PCA for dimension reduction. Table IV shows the MSE on plain audio generated without SSML attributes (baseline) and enhanced audio generated with SSML attributes predicted from text and character embeddings. Appending character embeddings to the input results in improvements over the baseline, but not over our best models.
@@ V-E Character Embeddings
TABLE IV: Test MSE for joint prediction of prosody attributes on dialogues using character embeddings (CE) different PCA dimensions.
VI Generating and Evaluating Text-to-Speech Audiobooks
We use the best performing model (LSTM with sequence length 3, and MPNet embeddings as input) to generate predictions for each book in the test set. We use a sliding window for inference, and compute the final prediction for each text input as the mean of its prediction in all windows. To convert the predicted z-scores into SSML attributes, we compute the mean and standard deviation for the pitch and volume in each human-read audio file in the test set, and use these values to convert the prosody attributes into absolute values (Hz for pitch and dB for volume). Using the mean pitch/volume as reference, we convert these into relative values as required by the SSML specification (as described in Section II-B). For rate of speech, we use a fixed value for mean and standard deviation (100 and 50 respectively) and perform the same computation.
We perform a human evaluation using Amazon Mechanical Turk. For each of the 24 books in the test set, we identify the phrases with the highest absolute values of predictions across the three prosody attributes. We then generate audio for this phrase text along with one context phrase before and after, using the Google Cloud Text-to-Speech API  [^7] https://cloud.google.com/text-to-speech with the voice en-US-Standard-A. We produce two different audio variants for each text:
@@ VI Generating and Evaluating Text-to-Speech Audiobooks
  • SSML-enhanced: We enclose each phrase in a <prosody> tag with the three predicted attributes specified: see example in Figure 2.
For each pair of audios, we ask five annotators to rate which audio sounds better and more natural with intonations similar to human readers. Figure 6 shows the results of the evaluation. For 12 out of 24 test sample pairs, a majority of the annotators preferred the SSML-enhanced version.
Figure 6: Distribution of books over annotator preference for SSML or Plain text.
VII Conclusion
We have developed an aligned text-to-audio book corpus, and deployed it to study both reader behavior and improve prosody models for TTS generation of audiobooks. We have shown that actors customize voices to represent specific characters, and adopt more expressive voices than those generated by conventional TTS systems, motivating the development of full-book analysis methods to understand/predict a reader’s goals in narrating a text.